OpenStudy (abbycross167):

Can someone please help me answer this question? I really need help ... I will fan and medal... Question~ write the following inequalities in y = mx + b form. 2x + y/3 < 6 *NOTE the < has a line under it

OpenStudy (abbycross167):

@retirEEd could you help me?

OpenStudy (retireed):

I hate inequalities ;( I normally treat them like they are NOT inequalities and add the inequality into the final answer. It is logical but it works for me.

OpenStudy (retireed):

Shall we start... write the following inequalities in y = mx + b form. 2x + y/3 < 6 2x + y/3 = 6 subtract 2x from both sides 2x -2x +y/3 = -2x +6

OpenStudy (retireed):

There is more, but what do you get after subtracting 2x from both sides.... 2x -2x +y/3 = -2x +6

OpenStudy (abbycross167):

y/3>6-2x?

OpenStudy (retireed):

> ???

OpenStudy (abbycross167):

sorry I meant to put =

OpenStudy (retireed):

Ok

OpenStudy (retireed):

So in order to get y/3= 6-2x Into this form y = mx + b What needs to happen?

OpenStudy (abbycross167):

Multiply both sides by 3?

OpenStudy (retireed):

Yes now do it.

OpenStudy (abbycross167):

3*y/3= 6-2x*3 y=6x+18

OpenStudy (retireed):

Close but not correct or maybe there is a typo.

OpenStudy (abbycross167):

hmmm I'm not sure is it a -6?

OpenStudy (retireed):

Yes -6x to be exact

OpenStudy (abbycross167):

oh ok

OpenStudy (retireed):

so now you need to add in the inequality

OpenStudy (abbycross167):

so -6+18=12?

OpenStudy (retireed):

What that make no sense to what we were trying to solve.

OpenStudy (abbycross167):

oh... I thought you meant to add those two numbers...

OpenStudy (retireed):

This was the original ... Question~ write the following inequalities in y = mx + b form. 2x + y/3 < 6

OpenStudy (retireed):

I meant the inequality sign. I shouldn't have said add, sorry.

OpenStudy (abbycross167):

oh ok, so Y>18-6x?

OpenStudy (retireed):

Nope

OpenStudy (abbycross167):

hmm :/ i don't know what to do then

OpenStudy (retireed):

Hint: 2x + y/3 < 6 *NOTE the < has a line under it

OpenStudy (abbycross167):

so I need to put the line under the < or would it be >?

OpenStudy (retireed):

Yes

OpenStudy (abbycross167):

ok umm do you know how to do that on a laptop?

OpenStudy (retireed):

Me either

OpenStudy (abbycross167):

ok I'll just add a note for my teacher... but other than the line this Y>18-6x is right?

OpenStudy (retireed):

but you need to make note of it Sometimes I do <= or >= Which one should it you use for this equation solution?

OpenStudy (abbycross167):

I think >

OpenStudy (abbycross167):

well =>

OpenStudy (retireed):

What was the original sign?

OpenStudy (abbycross167):

<= but I thought the sign changes?

OpenStudy (retireed):

Why would the sign change?

OpenStudy (abbycross167):

I don't know I just thought it did

OpenStudy (retireed):

The y is the dependent variable and x is the independent variable. Did the dependent variable change SIDEs ?

OpenStudy (abbycross167):

no sir/ma'am

OpenStudy (retireed):

So the inequality will not change. At least this is what I believe, but remember I don't like inequalities. Check with your teacher to be sure :)

OpenStudy (abbycross167):

yes sir, thank you very much!! you're amazing!

OpenStudy (retireed):

Thank you and ... y <= -6x +18 would be the proper form

OpenStudy (abbycross167):

I have one more question that I need help with... would you mind trying to help me it's graphing inequalities

OpenStudy (retireed):

Not tonight. I have to get ready for my internet talk show.

OpenStudy (abbycross167):

ok, well thank you for your help!